# LLAssist: Simple Tools for Automating Literature Review Using Large Language Models

### v2

@@ 2.2.1 Data Collection
    Scopus Database: using the search queries “llm AND cyber AND security” (term b), “llm OR ( generative AND artificial AND intelligence ) AND cyber AND security” (term c), and a broad “artificial AND intelligence AND cyber AND security” – limited to Conference paper and Article – (term d) to obtain a diverse dataset in a consecutively larger sample size.
@@ 4.1.2 Binary Relevance Decision and Score Distribution
The binary relevance decision and relevance score distribution obtained experiment are shown in Figure 2. Further, the summary of binary relevance and binary contribution decisions is shown in Table 1.
Figure 2: Binary Relevance and Relevance Store Distribution (small datasets)
@@ 4.2 Large Dataset Test
    Must-Read vs. Contributing Articles: While 324 articles (12.6%) are identified as must-read, only 100 articles (3.9%) are classified as potentially contributing. This suggests that LLAssist are more selective in identifying articles that directly contribute to answering the research questions, implying reduced time to manually read the abstract.

### v3

@@ 2.2.1 Data Collection
    Scopus Database: using the search queries “llm AND cyber AND security” (term b), “llm OR ( generative AND artificial AND intelligence ) AND cyber AND security” (term c), and a broad “artificial AND intelligence AND cyber AND security” – limited to Conference paper – (term d) to obtain a diverse dataset in a consecutively larger sample size.
@@ 4.1.2 Binary Relevance Decision and Score Distribution
The binary relevance decision and relevance score distribution obtained from the experiment are shown in Figure 2. Further, the summary of binary relevance and binary contribution decisions is shown in Table 1.
Figure 2: Binary Relevance and Relevance Score Distribution (small datasets)
@@ 4.2 Large Dataset Test
    Must-Read vs. Contributing Articles: While 324 articles (12.6%) are identified as must-read, only 100 articles (3.9%) are classified as potentially contributing. This suggests that LLAssist is more selective in identifying articles that directly contribute to answering the research questions, implying reduced time to manually read the abstract.